# Model predictive control is almost optimal for restless bandits

### v1

@@ Abstract
We consider the discrete time infinite horizon average reward restless markovian bandit (RMAB) problem. We propose a model predictive control based non-stationary policy with a rolling computational horizon $\tau$. At each time-slot, this policy solves a $\tau$ horizon linear program whose first control value is kept as a control for the RMAB. Our solution requires minimal assumptions and quantifies the loss in optimality in terms of $\tau$ and the number of arms, $N$. We show that its sub-optimality gap is $O(1/\sqrt{N})$ in general, and $\exp(-\Omega(N))$ under a local-stability condition. Our proof is based on a framework from dynamic control known as dissipativity. Our solution is easy to implement and performs very well in practice when compared to the state of the art. Further, both our solution and our proof methodology can easily be generalized to more general constrained MDP settings and should thus, be of great interest to the burgeoning RMAB community.
@@ 1 Introduction
This work investigates the sequential decision making problem of Restless Multi-Armed Bandits (RMAB for short) over an infinite discrete time horizon. In this problem there are $N$ statistically identical arms. At each time step, the decision maker must choose for each arm if they would like to pull the arm or leave it as is. The decision maker has a constraint $\alpha N$ on the maximal number of arms that they may pull at each time instance. Each arm has a known state belonging to a common finite state space and upon choosing an action, produces a known state and action dependent reward. Next, the arms evolve to a new state independently according to a known state-action dependent transition kernel. These arms are only coupled through the budget constraint on the number of arms that maybe pulled at each time instance. The state and reward are both revealed to the decision maker before the next decision needs to be made. The objective of the decision maker is to maximize the long term time average reward.
@@
This problem was first proposed by Whittle (1988). Over the years RMABs have been used to model a number of practical problems. These applications include web-crawling, queuing, communication systems, scheduling problems and many more, (Veatch & Wein, 1996), (Dance & Silander, 2019), (Nino-Mora, 2002), (O’Meara & Patel, 2001). The problem of choosing a subset of tasks to perform among a larger collection of tasks under resource constraints shows up time and time again in various resource constrained control problems. For a comprehensive review on RMABs and their applications, the interested reader is directed towards Niño-Mora (2023). While the existence of an optimal policy for RMABs straightforward, Papadimitriou & Tsitsiklis (1999) showed that the exact solution to this problem is PSPACE-hard. Consequently, most work focused on designing approximate solutions with good performance guarantees.
@@
In the seminal paper, Whittle (1988) suggested that under a condition known as indexability, an index can be associated with each state. This index is now referred to as the Whittle’s index (WI) and it was conjectured that a priority policy based on this index would be an optimal solution for this problem. This setting naturally lends itself to mean field approximations where one may replace the $N$ armed problem with a dynamical system in order to find these approximate solutions. (Weber & Weiss, 1990) were the first to point out that under a (hard to verify) condition on the dynamical system known as uniform global attractor property (UGAP), the (WI) was asymptotically optimal. Recently, many of the results in the RMAB literature have focused on two major aspects of the problem. Firstly, how quickly do proposed asymptotically optimal policies converge to the optimal solution as a function of the number of arms. Secondly, can the underlying assumptions such as indexability and UGAP be made less restrictive to obtain more general conditions under which optimal solutions can be found. In the former category, under the assumptions of Weber & Weiss (1990), Gast et al. (2023a) showed that the WI policy is exponentially close to the optimal solution. More recently still, several works have been able to show an exponential order of convergence, Gast et al. (2023b); Hong et al. (2024a) under less restrictive assumptions. On the other hand, in the latter category, Hong et al. (2023) showed an asymptotic convergence result under a far less restrictive assumption known as the synchronization assumption. Similar works following this steering dynamics include Yan (2024) and Hong et al. (2024b) to show asymptotically optimal algorithms. Furthermore, Hong et al. (2024a) loosened the restrictions further to not only show asymptotic convergence but describe fundamental conditions in order to achieve exponential convergence rates for any algorithm. A more comprehensive view of recent results can be found in Appendix A with descriptions of both assumptions and the algorithm.
In this work, we return to the dynamical control perspective but do not look to steer our system towards an optimal fixed point. Instead, we focus on the notion of dissipativity, to make the case for a well known set of policies known as model predictive control. We show that a simple MPC with a finite planning horizon produces an asymptotically optimal solution to the RMAB problem without any need to relax constraints. In doing so, not only do we bring a new perspective to this decades old problem but also describe a simple tractable algorithm that performs exceptionally well in practice with a small planning horizon.
@@
Our main contribution through this paper is to show that a simple computationally tractable algorithm with minimal assumptions is asymptotically optimal in the number of arms. Along the way we make several fresh observations about the nature of RMAB problems:
  1. 1.
    First and foremost, we show that the sub-optimality gap of the natural MPC algorithm (called LP-update in the literature) is $\mathcal{O}(1/\sqrt{N})$, where $N$ is the number of arms. This result uses a minimal set of assumptions.
  2. 2.
@@
  4. 4.
    Perhaps the most helpful portion of our results, for practical purposes; the MPC algorithm works well in practice and is easy to implement. It performs well both in terms of the number of arms $N$ as well as the computational time horizon $T$ beating state of the art algorithms in our bench marks.
Roadmap
The rest of the paper is organized as follows. We describe the system model and the corresponding linear relaxation in Section 2. We build the LP-update algorithm in Section 3 and present its performance guarantee in Section 4. We provide the main ingredients of the proof in Section 5 postponing the most technical lemmas to the appendix. We illustrate the performance of the algorithm in Section 6. The appendix contains additional literature review A, details of the algorithm and their extension to multi-constraints MDPs B, additional proofs C, D and details about the numerical experiments F.
@@ 2.1 System Model
We consider an infinite horizon discrete time restless Markovian bandit problem parameterized by the tuple $\langle\mathcal{S},\mathbf{P}^{0},\mathbf{P}^{1},\mathbf{r}^{0},\mathbf{r}^{1};\alpha,N\rangle$. A decision maker is facing $N$ statistically identical arms, and each arm has a state that belongs to the finite state-space $\mathcal{S}$. At each time-instant, the decision maker observes the states of all arms $\textbf{s}=\{s_{1}\dots s_{N}\}$ and chooses a vector of actions $\textbf{a}=\{a_{1}\dots a_{N}\}$, where the action $a_{n}=1$ (respectively $a_{n}=0$) corresponds to pulling the arm $n$ (or to leave it). The decision maker is constrained to pull at most $\alpha N$ arms at each decision epoch.
The matrices $\mathbf{P}^{0}$ and $\mathbf{P}^{1}$ denote the transitions matrices of each arm and the vectors $\mathbf{r}^{0},\mathbf{r}^{1}$ denote the $|\mathcal{S}|$ dimensional vector for the rewards. We assume that all the rewards $\mathbf{r}$ lie between 00 and $1$. As the state-space is finite, this assumption can be made without loss of generality by scaling and centering the reward vector. We assume that the transitions of all arms are Markovian and independent. This means that if the arms are in state $\textbf{s}\in\mathcal{S}^{N}$ and the decision maker takes an action $\textbf{a}\in\{0,1\}^{N}$, then the decision maker earns a reward $\sum_{n}r^{n}_{a_{n}}$ and the next state becomes $\textbf{s}^{\prime}\in\mathcal{S}^{N}$ with probability
@@
Let $\pi$ be a stationary policy mapping each state to a probability of choosing actions, i.e, $\pi:\mathcal{S}^{N}\to\Delta(\{0,1\}^{N})$ subject to the budget constraint, $\alpha$. Let $\Pi^{(N)}$ denote the space of all such policies. Given a policy $\pi$, mapping the joint state of the arms to a realization of the $N$ length joint action vector at time $t$, $\textbf{A}^{N}(t):=\{a_{1}(t),a_{2}(t)\dots a_{N}(t)\}$, and an initial set of arm states $\textbf{S}^{(N)}(0)=\textbf{s}$, we define the average gain of policy $\pi$ as
$V_{\pi}^{(N)}(\textbf{s})=\lim_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\frac{1}{N}\sum_{n=1}^{N}r^{A_{n}(t)}_{S_{n}(t)}\Bigg{|}\textbf{S}^{(N)}(0)=\textbf{s}\right].$ (2)
Here $(S^{N}_{n}(t),A^{N}_{n}(t))$ denotes the state-action pair of the $n$th arm at time $t$ and $r^{A^{N}_{n}(t)}_{S^{N}_{n}(t)}$ denotes the $S^{N}_{n}(t)^{\text{th}}$ entry of the $\mathbf{r}^{A_{n}(t)}$ vector. As the state-space and action space is finite, for any stationary policy $\pi$, the limit is well defined Puterman (2014). The RMAB problem amounts to computing a policy $\pi$ that maximizes the infinite average reward. We denote the optimal value of the problem as
$V_{\text{opt}}^{N}(\textbf{s})$ $:=\max_{\pi\in\Pi^{(N)}}V_{\pi}^{(N)}(\textbf{s}).$ (3)
. The optimal policy exists and the limit is well defined, Puterman (2014). Under mild conditions (which will be verified in our case), this value does not depend on the initial state in which case we will simply write it as $V_{\text{opt}}^{N}$.
@@ 2.3 Linear relaxation
In (5), the variable $x_{i}$ corresponds to the time-averaged fraction of arms in state $i$ and similarly the variable $u_{i}$ corresponds to the time-averaged fraction of arms in state $i$ that are pulled. The constraint (5b) imposes that on average, no more than $\alpha N$ arms are pulled. This is in contrast with the condition imposed for problem (3) that enforces this condition at each time step.
@@ 3.1 The Finite-Horizon Mean Field Control Problem
Before moving forward, the above equation deserves some remarks. First, for any finite $\tau$, the objective and the constraints (7) of the optimization problem (8) are linear in the variables $({\bm{x}}(t),\bm{u}(t))$. This means that this optimization problem is computationally easy to solve. In what follows, we denote by $\mu_{\tau}(\textbf{x})$ the value of $\bm{u}(0)$ of an optimal solution to (8).
Second, the definition of (8) imposes that the constraint $\|\bm{u}\|_{1}\leq\alpha\|\textbf{x}\|_{1}=\alpha$ holds for each time $t$. This is in contrast to the way this constraint is typically treated in RMAB problems, in which case (4) is replaced with the time-averaged constraint $\frac{1}{T}\sum_{t=0}^{T-1}\|\bm{u}(t)\|_{1}\leq\alpha$. The latter relaxation was introduced in Whittle (1988) and is often referred to as Whittle’s relaxation (Avrachenkov & Borkar, 2020; Avrachenkov et al., 2021). This is the constraint that we use to write (5). Gast et al. (2023a) showed that for any finite $T$, the finite-horizon equivalent of (3) converges to (8) as $N$ goes to infinity. The purpose of this paper is to show that the solution of the finite $T-$horizon LP (8) provides an almost-optimal solution to the original $N-$arm average reward problem (3).
Last, as we will discuss later, taking $\lambda$ as the dual multiplier of the constraint (5c) helps to make a connection between the finite and the infinite-horizon problems (8) and (5). Our proofs will hold with minor modification by replacing $\lambda$ by 00 and in practise we do not use this multiplier for our algorithm.
@@ 3.2 The Model Predictive Control Algorithm
  Average $\text{Average reward}:\frac{\text{Total-reward}}{T}\approx V_{\text{LP},\tau}^{(N)}(\infty)$
The pseudo-code of the LP-update policy is presented in Algorithm 1. The LP-update policy takes as an input a time-horizon $\tau$. At each time-slot, the policy solves the finite horizon linear program (8) to obtain $\mu_{\tau}(\textbf{x})$ that is the value of $\bm{u}(0)$ of an optimal solution to (8). Note that such a policy may not immediately translate to an applicable policy as we do not require that $N\mu_{\tau}(\textbf{x})$ to be integers. We therefore use randomized rounding to obtain a feasible policy for our $N$ armed problem, $\textbf{A}^{N}(t)$. Applying these actions to each arm gives an instantaneous reward and a next state. This form of control has been referred to as rolling horizon Puterman (2014) but more commonly referred to as model predictive control Damm et al. (2014). Our algorithm maybe summarized more succinctly as:
@@
The randomized rounding procedure that we use is similar to the one described in Gast et al. (2023a). We discuss it in Appendix B.1.
@@ 4 Main theoretical results
The main result of our paper is to show that a finite-horizon model predictive control algorithm, that we call the LP-update policy, is asymptotically optimal for the infinite horizon bandit problem. Note that this LP-update policy is introduced in Gast et al. (2023a; b); Ghosh et al. (2022) for finite-horizon restless bandit.

### v2

@@ Abstract
We consider the discrete time infinite horizon average reward restless markovian bandit (RMAB) problem. We propose a model predictive control based non-stationary policy with a rolling computational horizon $\tau$. At each time-slot, this policy solves a $\tau$ horizon linear program whose first control value is kept as a control for the RMAB. Our solution requires minimal assumptions and quantifies the loss in optimality in terms of $\tau$ and the number of arms, $N$. We show that its sub-optimality gap is $O(1/\sqrt{N})$ in general, and $\exp(-\Omega(N))$ under a local-stability condition. Our proof is based on a framework from dynamic control known as dissipativity. Our solution is easy to implement and performs very well in practice when compared to the state of the art. Further, both our solution and our proof methodology can easily be generalized to more general constrained MDP settings and should thus be of great interest to the burgeoning RMAB community.
keywords:
Restless Bandits, LP-update policies, Constrained MDPs
@@ 1 Introduction
This work investigates the sequential decision making problem of Restless Multi-Armed Bandits (RMAB for short) over an infinite discrete time horizon. In this problem there are $N$ statistically identical arms. At each time step, the decision maker must choose for each arm if they would like to pull the arm or leave it as is. The decision maker has a constraint $\alpha N$ on the maximal number of arms that they may pull at each time instance. Each arm has a known state belonging to a common finite state space and upon choosing an action, produces a known state and action dependent reward. Next, the arms evolve to a new state independently according to a known state-action dependent transition kernel. These arms are only coupled through the budget constraint on the number of arms that may be pulled at each time instance. The state and reward are both revealed to the decision maker before the next decision needs to be made. The objective of the decision maker is to maximize the long-term time average reward.
@@
This problem was first proposed by Whittle (1988). Over the years RMABs have been used to model a number of practical problems. These applications include web-crawling, queuing, communication systems, scheduling problems and many more, (Veatch and Wein, 1996), (Dance and Silander, 2019), (Nino-Mora, 2002), (O’Meara and Patel, 2001). The problem of choosing a subset of tasks to perform among a larger collection of tasks under resource constraints shows up time and time again in various resource constrained control problems. For a comprehensive review on RMABs and their applications, the interested reader is directed towards Niño-Mora (2023). While the existence of an optimal policy for RMABs is straightforward, Papadimitriou and Tsitsiklis (1999) showed that the exact solution to this problem is PSPACE-hard. Consequently, most work focused on designing approximate solutions with good performance guarantees.
@@
In the seminal paper, Whittle (1988) suggested that under a condition known as indexability, an index can be associated with each state. This index is now referred to as the Whittle’s Index (WI), and it was conjectured that a priority policy based on this index would be an optimal solution for this problem. This setting naturally lends itself to mean field approximations where one may replace the $N$ armed problem with a dynamical system in order to find these approximate solutions. (Weber and Weiss, 1990) were the first to point out that under a (hard to verify) condition on the dynamical system known as uniform global attractor property (UGAP), the (WI) was asymptotically optimal. Recently, many of the results in the RMAB literature have focused on two major aspects of the problem: how quickly do proposed asymptotically optimal policies converge to the optimal solution as a function of the number of arms and whether the underlying assumptions such as indexability and UGAP can be made less restrictive to obtain more general conditions under which optimal solutions can be found. In the former category, under the assumptions of Weber and Weiss (1990), Gast et al. (2023a) showed that the WI policy is exponentially close to the optimal solution. Recently, several works have been able to show an exponential order of convergence, Gast et al. (2023b); Hong et al. (2024a) without an indexability condition. On the other hand, in the latter category, Hong et al. (2023) showed an asymptotic convergence result under a less restrictive assumption known as the synchronization assumption. Similar works include Yan (2024) and Hong et al. (2024b) which show asymptotically optimal algorithms. Furthermore, Hong et al. (2024a) loosen the restrictions to show asymptotic convergence and describe fundamental conditions to achieve exponential convergence rates for any algorithm. Our paper falls partially in the latter category by providing asymptotic convergence results under the weakest assumptions (to the best of our knowledge) but we also show that under certain local stability assumptions one retrieves exponential convergence rates. A more comprehensive view of recent results can be found in the supplementary material.
@@
The main contribution of our paper is the proof that a very natural model predictive control (LP-update) provides the best of both worlds, it not only provides an algorithm that works well in practice, but we can also theoretically guarantee under the weakest assumptions (to the best of our knowledge) convergence to the optimal policy. This idea of resolving an LP for a finite-horizon has been used previously; however, its use has been entirely restricted to either the finite-horizon case or, when used in the infinite horizon case, required a critical UGAP assumption(uniform global attractor property) on the system. This assumption is known to be extremely hard to verify Gast et al. (2023a). In this paper, by introducing the framework of dissipativity we are able to show that a finite horizon policy indeed provides an almost optimal policy for the infinite horizon problem. Note, in the absence of such a framework, the finite horizon policy may veer very far from the infinite horizon policy Damm et al. (2014). The use of this framework is one of the key technical novelties of our approach. Apart from the main result, we make several fresh observations about the nature of RMAB problems:
@@
  • Perhaps the most helpful portion of our results, for practical purposes: the MPC algorithm works well in practice and is easy to implement. It performs well both in terms of the number of arms $N$ as well as the computational time horizon $T$, beating state-of-the-art algorithms in our bench marks.
  • Finally, from a control systems point of view, we return from trying to steer a dynamical system towards a fixed point to trying to maximize a value function. Policies can often be discontinuous in states while value functions are often smoother under mild assumptions.
Road-map
The rest of the paper is organized as follows. We describe the system model and the corresponding linear relaxation in Section 2. We build the LP-update algorithm in Section 3 and present its performance guarantee in Section 4. We provide the main ingredients of the proof in Section 5 postponing the most technical lemmas to the appendix. We illustrate the performance of the algorithm in Section 6. The appendix contains additional literature review A, details of the algorithm and their extension to multi-constraints MDPs B, additional proofs C, D and details about the numerical experiments G.
Reproducibility
The code to reproduce this paper (including all simulations and figures) is available at https://github.com/ngast/paper_banditMPC_colt2025.
@@ 2.1 System Model
We consider an infinite horizon discrete time restless Markovian bandit problem parameterized by the tuple $\langle\mathcal{S},\mathbf{P}^{0},\mathbf{P}^{1},\mathbf{r}^{0},\mathbf{r}^{1};\alpha,N\rangle$. A decision maker is facing $N$ statistically identical arms, and each arm has a state that belongs to the finite state-space $\mathcal{S}$. At each time-instant, the decision maker observes the states of all arms $\textbf{s}=\{s_{1}\dots s_{N}\}$ and chooses a vector of actions $\textbf{a}=\{a_{1}\dots a_{N}\}$, where the action $a_{n}=1$ (respectively $a_{n}=0$) corresponds to pulling the arm $n$ (or leaving it). The decision maker is constrained to pull at most $\alpha N$ arms at each decision epoch.
The matrices $\mathbf{P}^{0}$ and $\mathbf{P}^{1}$ denote the transitions matrices of each arm and the vectors $\mathbf{r}^{0},\mathbf{r}^{1}$ denote the $|\mathcal{S}|$ dimensional vector for the rewards. We assume that all the rewards $\mathbf{r}$ lie between 00 and $1$. As the state-space is finite, this assumption can be made without loss of generality by scaling and centering the reward vector. We assume that the transitions of all arms are Markovian and independent. This means that if the arms are in state $\textbf{s}:=\{s_{1},s_{2}\dots s_{N}\}\in\mathcal{S}^{N}$ and the decision maker takes an action $\textbf{a}\in\{0,1\}^{N}$, then the decision maker earns a reward $\frac{1}{N}\sum_{n}r^{a_{n}}_{s_{n}}$ and the next state becomes $\textbf{s}^{\prime}:=\{s^{\prime}_{1},s^{\prime}_{2}\dots s^{\prime}_{N}\}\in\mathcal{S}^{N}$ with probability
@@
Let $\pi$ be a stationary policy mapping each state to a probability of choosing actions, i.e, $\pi:\mathcal{S}^{N}\to\Delta(\{0,1\}^{N})$ subject to the budget constraint $\alpha$. Let $\Pi^{(N)}$ denote the space of all such policies. We define the average gain of policy $\pi\in\Pi^{(N)}$ as
$V_{\pi}^{(N)}=\lim_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\frac{1}{N}\sum_{n=1}^{N}r^{A_{n}(t)}_{S_{n}(t)}\right].$ (2)
In theory, the average gain of a policy might depend on the initial state $\textbf{S}^{(N)}(0)$. Yet, under mild conditions (which will be verified in our case), this value does not depend on the initial state. This is why we omit the dependence on the initial state and simply write $V_{\pi}^{(N)}$.
Here $(S_{n}(t),A_{n}(t))$ denotes the state-action pair of the $n$th arm at time $t$ and $r^{A_{n}(t)}_{S_{n}(t)}$ denotes the $S_{n}(t)^{\text{th}}$ entry of the $\mathbf{r}^{A_{n}(t)}$ vector. As the state-space and action space is finite, for any stationary policy $\pi$, the limit is well defined, Puterman (2014). The RMAB problem amounts to computing a policy $\pi$ that maximizes the infinite average reward. We denote the optimal value of the problem as
$V_{\text{opt}}^{N}$ $:=\max_{\pi\in\Pi^{(N)}}V_{\pi}^{(N)}.$ (3)
The optimal policy exists and the limit is well defined, Puterman (2014).
@@ 2.3 Linear relaxation
In (5), the variable $x_{i}$ corresponds to the time-averaged fraction of arms in state $i$; similarly the variable $u_{i}$ corresponds to the time-averaged fraction of arms in state $i$ that are pulled. The constraint (5b) imposes that on average, no more than $\alpha N$ arms are pulled. This is in contrast with the condition imposed for problem (3) that enforces this condition at each time step.
@@ 3.1 The Finite-Horizon Mean Field Control Problem
Before moving forward, the equation above deserves some remarks. First, for any finite $\tau$, the objective and the constraints (7) for the optimization problem (8) are linear in the variables $({\bm{x}}(t),\bm{u}(t))$. This means that this optimization problem is computationally easy to solve. In what follows, we denote by $\mu_{\tau}(\textbf{x})$ the value of $\bm{u}(0)$ of an optimal solution to (8).
Second, the definition of (8) imposes that the constraint $\|\bm{u}\|_{1}\leq\alpha\|\textbf{x}\|_{1}=\alpha$ holds for each time $t$. This is in contrast to the way this constraint is typically treated in RMAB problems, in which case (4) is replaced with the time-averaged constraint $\frac{1}{T}\sum_{t=0}^{T-1}\|\bm{u}(t)\|_{1}\leq\alpha$. The latter relaxation was introduced in Whittle (1988) and is often referred to as Whittle’s relaxation (Avrachenkov and Borkar, 2020; Avrachenkov et al., 2021). This is the constraint that we use to write (5). Gast et al. (2023a) showed that for any finite $T$, the finite-horizon equivalent of (3) converges to (8) as $N$ goes to infinity. The purpose of this paper is to show that the solution of the finite $T-$horizon LP (8) provides an almost-optimal solution to the original $N-$arm average reward problem (3).
Last, as we will discuss later, taking $\lambda$ as the dual multiplier of the constraint (5c) helps to make a connection between the finite and the infinite-horizon problems (8) and (5). Our proofs will hold with minor modification by replacing $\lambda$ by 00 and in practice we do not use this multiplier at all.
@@ 3.2 The Model Predictive Control Algorithm
  Average $\text{Average reward}:\frac{\text{Total-reward}}{T}\approx V_{\text{LP -}\tau}^{(N)}$
The pseudo-code of the LP-update policy is presented in Algorithm 1. The LP-update policy takes as an input a time-horizon $\tau$. At each time-slot, the policy solves the finite horizon linear program (8) to obtain $\mu_{\tau}(\textbf{x})$ that is the value of $\bm{u}(0)$ for an optimal solution to (8). Note that such a policy may not immediately translate to an applicable policy as we do not require that $N\mu_{\tau}(\textbf{x})$ be integers. We therefore use randomized rounding to obtain a feasible policy for our $N$ armed problem, $\textbf{A}^{N}(t)$. Applying these actions to each arm gives an instantaneous reward and a next state. This form of control has been referred to as rolling horizon Puterman (2014) but is more commonly referred to as model predictive control Damm et al. (2014). Our algorithm may be summarized by:
@@
We use a randomized rounding procedure similar to Gast et al. (2023a), see Appendix B.1.
@@ 4 Main theoretical results
The main result of our paper is to show that a finite-horizon model predictive control algorithm, that we call the LP-update policy, is asymptotically optimal for the infinite horizon bandit problem. Note that this LP-update policy is introduced in Gast et al. (2023a, b); Ghosh et al. (2022) for finite-horizon restless bandit.